Find the pH of 0.25 M HF(aq) given that its Ka = 3.5×10-4.

3.46

4.06

0.60

2.03

1.73
Answer #1

H+ + - HE 0.25 - X 0.25 x Ka= [H + ] [f-1 HF 3.5x10 Y= x2 0.25-2 x² + 3.5*10tk- 8.75x10-sol solving for a x = 9.18x163 H+ 1 =